solve for equation for; four teimes the sum of a number and three is equal to eight more than twice the same number?
Found 2 solutions by stanbon, sachi:
Answer by stanbon(75887)

solve for equation for; four teimes the sum of a number and three is equal to eight more than twice the same number?
-------
Equation:
4(x + 3) = 2x+8
----------------------
4x + 12 = 2x + 8
------
2x = -4
x = -2
=============
Cheers,
Stan H.
=============

Answer by sachi(548)

let the no is x
four times the sum of a number and three is equal to eight more than twice the same number
so 4(x+3)=8+2x
or 4x+12=8+2x
or 4x-2x=8-12=-4
or 2x=-4
or x=-2